OpenStudy (anonymous):

what is 4 with exponent -1/2

OpenStudy (anonymous):

\[4^{(-1/2)}\]

OpenStudy (anonymous):

the negative in the exponent means that the expression can be in this form too: \[1/(4^{1/2})\]

OpenStudy (anonymous):

1/8

OpenStudy (anonymous):

so, you need to calculate 4^1/2 which the square root of 4

OpenStudy (anonymous):

1/2

OpenStudy (anonymous):

my bad

OpenStudy (anonymous):

sorry but what is the ^ mean?

OpenStudy (anonymous):

that is represents an exponential expression...so 4^(1/2).... is \[4^{1/2}\]

OpenStudy (anonymous):

which is related to multiplication and division

OpenStudy (anonymous):

the "^" is used when typing because it's usually difficult to do a superscript

OpenStudy (anonymous):

thx the carrot info makes sense. Unfortunately the exponent answer does not :(

OpenStudy (anonymous):

ok, what about it?

OpenStudy (anonymous):

4^-1/2 = ?

OpenStudy (radar):

\[4^{-1/2}=1/4^{1/2}=1/\sqrt{4}=\pm1/2\]

OpenStudy (anonymous):

wow.

OpenStudy (anonymous):

um..........4 1/2..........???????

OpenStudy (anonymous):

nova, a fractional exponent means you're taking a root. an exponent of 1/2 means a square root. an exponent of 1/3 means a cubed root.

OpenStudy (anonymous):

sammy, 4 to the 1/2 power is same as square root of 4.

OpenStudy (anonymous):

oh............remember i'm only 10

OpenStudy (anonymous):

sweet! welcome aboard!

OpenStudy (anonymous):

The "^" symbol they used is shorthand for "exponents"

OpenStudy (anonymous):

thx

OpenStudy (anonymous):

i am a fan of Nauq

OpenStudy (anonymous):

Thanks Sammy.

OpenStudy (anonymous):

your welcome. lets group chat